OpenStudy (anonymous):

Solve And Check. 16=1/4p A. 1/4 B. 4 C. 20 D. 64 please help!!!!

rvc (rvc):

cross multiply

OpenStudy (anonymous):

D

OpenStudy (anonymous):

(4)1/4p=16(4) p=16(4) p=64

OpenStudy (anonymous):

Ok, check this website out: http://www.tiger-algebra.com/drill/16=1/4p/ It shows you how to solve this problem step-by-step ;)

OpenStudy (anonymous):

NP :)